# LASLiN: A Learning-Augmented Peer-to-Peer Network

### v1

@@ 3 Optimum Static Skip List Network (OSSLN)
OSSLN with Bounded Height: In the OSSLN with bounded height, the height assignment is of the form $h:V\longrightarrow[1,h_{\max}]$, where $h_{\max}$ is a parameter, typically bounded by $O(\log{n})$ or $O(1)$.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$. Let $nl$, $1\leq nl\leq l$, denote the leftmost vertex id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost vertex id, such that $\forall v\in[r,nr],\,h_{v}\leq h$. Finally, let $\text{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$.
Theorem 3.3.
If $x\in[l,r]$ is a candidate pivot vertex with height $h\in[1,h_{\max}]$, then the following recurrence relation holds $\forall(nl,l,x,r,nr)$, $1\leq nl\leq l\leq x\leq r\leq nr\leq n$:
$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
Proof 3.4.
Base case $(l=r)$: In this case, we have that $(x=l)$ and the set of paths $x$ belongs to is given by Lemma 3.1. By setting $nl=L(x,h)$ and $nr=R(x,h)$ in the proof, the vertex contribution to the total cost is given by:
Let $1\leq l<r\leq n$ and $h\in[1,h_{\max}]$. Assume that the recurrence equation is true for all $h^{\prime},l^{\prime},r^{\prime}$ such that $(h^{\prime}<h)\leavevmode\nobreak\ \wedge\leavevmode\nobreak\ (r^{\prime}-l^{\prime})<(r-l)$. From previous step also assume that the equation is true for $(r^{\prime}-l^{\prime})=(r-l)$ with height $(h^{\prime}<h)$.
$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
and OSSLN(l,r,l,r,h1,W)OSSLN(l,r,l,r,h-1,W), which is precisely what we want to prove.
If we pre-compute an auxiliary matrix of aggregate weights ($W^{\prime}(x,y)=\sum_{z\leq y}W(x,z),\forall(x,y),1\leq x\leq y\leq n$) in time $O(n^{3})$, then the above dynamic programming algorithm has complexity $\mathcal{O}(n^{5}\cdot h_{\max})$.
OSSLN with Unbounded height (OSSLNU): We now present a dynamic programming algorithm for the OSSLN problem with unbounded heights, i.e., when the height assignment is of the form $h:V\longrightarrow[1,n]$. We show that, in this particular case, the algorithm’s complexity can be reduced from $\mathcal{O}(n^{5}\cdot h_{\max})$ to $\mathcal{O}(n^{3})$.
A crucial observation is that, for any solution to the OSSLNU problem, there exists a solution of cost less than or equal to that of the former, in which, $\forall[l,r],1\leq l\leq r\leq n$, there is a unique pivot vertex, i.e., a unique vertex whose height is the maximum height in that interval.
Lemma 3.5.
Let $h$ be an optimum solution to $OSSLNU(V,W)$. Then, there must exist a solution $h^{\prime}$ such that $Cost((V,h^{\prime},E^{\prime}),W)\leq Cost((V,h,E),W)$, where, for any interval $l,r],1\leq l\leq r\leq n,$ there is exactly one vertex $x\in[l,r]$ with height $h^{\prime}_{x}=\max_{w\in[l,r]}h^{\prime}_{w}$.
Proof 3.6.
@@
Since $\text{Cost}((V,h,E),W)=\sum_{(u,v)\in V}W(u,v)\cdot d(u,v)$ and $d(u,v)=|P_{h}(u,v)|$, and any path in $h^{\prime}$ is of at most the same size as in $h$, then, $\text{Cost}((V,h^{\prime},E^{\prime}),W)\leq\text{Cost}((V,h,E),W)$. Hence, the lemma holds.
In OSSLNU, because we are only interested in the relative order of vertex heights, rather than in their absolute values, we have that $(nl=l)$ and $(nr=r)$, $\forall[l,r],1\leq l\leq r\leq n$, and the height parameter $h$ becomes unnecessary. This greatly simplifies the recurrence relation, reducing the complexity to $\mathcal{O}(n^{3})$.
Theorem 3.7.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$. Let $\text{OSSLNU}(l,r,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$. Let $x\in[l,r]$ be a candidate pivot vertex. Define $l$ as the leftmost index such that $\forall x,\in[l,r],\forall v\in[l,x],\,h_{v}\leq h_{x}$, and similarly define $r$ as the rightmost index such that $\forall v\in[x,r],\,h_{v}\leq h_{x}$. Then:
$\text{OSSNU}(l,r,W)=\min_{x\in[l,r]}\Big{(}\text{OSSNU}(l,x-1,W)+$
$\text{OSSNU}(x+1,r,W)+$
$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)\Big{)}.$ (5)
Proof 3.8.
@@
Note that, by Lemma 3.5, once a pivot vertex $x\in[l,r]$ is chosen as the highest vertex within the interval $[l,r]$, we can assume, without loss of generality, that all vertices in the optimal solutions to the left subinterval [l,x1][l,x-1] and to the right subinterval $[x+1,r]$ have heights strictly lower than $h_{x}$.
@@
Base Case $(l=r)$: In this case, the interval contains a single vertex $x=l=r$. In this case, the cost reduces to the contribution of $x$ to the solution cost, which, from Lemma 3.1, and $L(x,h)=l$ and $R(x,h)=r$ as we demonstrated above, it follows that this expression corresponds exactly to:
@@
  • The cost of all paths that include $x$, which by Lemma 3.1 is given by the Eq.(6).
Since we minimize this total over all pivot choices $x\in[l,r]$, the recurrence holds.
4 Continuous SLN
This section introduces Continuous SLN, a generalization of SLN.
Definition 4.1 (Continuous SLN).
A Continuous Skip List Network (Continuous SLN) $\mathcal{N}=(V,h,E)$ is an SLN (Definition 2.1) except the height function $h$ is defined over the real numbers, i.e., $h:V\longrightarrow\mathbb{R}$.
We prove that continuous skip graphs support a functioning P2P network by implementing the three primitives: route, join and leave.
Internal Variables.
A node $u$ holds the variables:
@@ 4 Continuous SLN
The route algorithm (Algorithm 1) takes the destination node id $v$ and hops from node to node until it finds the destination node, returning $(v,True)$ or reaches a dead end in some node $w\neq v$, return $(w,False)$.
1 // code running at node $u$
2 if $u$ = $v$ then
3   return ($u$, True)
4   else if $u<v$ then
5     $S\leftarrow\{x\in N_{u}|u<x\leq v\}$
6     if $S=\emptyset$ then
7       return ($u$, False)
8       else
9         $w\leftarrow\max S$
10         return $w.\text{route}(v)$
11
12         else
13           $S\leftarrow\{x\in N_{u}|v\leq x<u\}$
14           if $S=\emptyset$ then
15             return ($u$, False)
16             else
17               $w\leftarrow\min S$
18               return $w.\text{route}(v)$
19
Algorithm 1 route($v$)
@@ 17               $w\leftarrow\min S$
The join algorithm (Algorithm 2) inserts the node inside the Continuous SLN based on its anchor local variable. The potential neighbor changes of the other nodes propagate through the network with the find_neighbors algorithm (Algorithm 3).
1
2 // code running at node $u$
3 if $\text{anchor}<u$ then
4   $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
5
6   else
7     $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
8
9     if $S=\emptyset$ then
10       $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
11       Update $D_{u}$
12       else
13         if $\text{anchor}<u$ then
14           $\text{succ}\leftarrow\min S$
15
16           else
17             $\text{succ}\leftarrow\max S$
18
19             $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
20             $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{find\_neighbors}(u,h_{v},-\infty)$
21             Update $D_{u}$
22
Algorithm 2 join()
1
2 // code running at node $u$
3 if current_height = -$\infty$ then
4   $\text{anchor}_{u}\leftarrow v$
5
6  $N\leftarrow\emptyset$
7   if current_height $h_{u}$ then
8     $N\leftarrow\{u\}$
9     $N_{u}\leftarrow N_{u}\cup\{v\}$
10     Update $D_{u}$
11     if $u<v$ then
12       $N\leftarrow\{x\in N_{u}|x\leq v\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
13       for $x\in N_{u}$ do
14         if $x<u\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
15           $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
16
17
18
19          else
20             $N\leftarrow\{x\in N_{u}|v\leq x\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
21             for $x\in N_{u}$ do
22               if $u<x\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
23                 $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
24
25
26                 return NN
Algorithm 3 find_neighbors($v$, $h_{v}$, current_height)
@@ 10     Update $D_{u}$
The leave algorithm (Algorithm 4, in this paper’s appendix) shows how an already-inserted node can leave the network. The potential neighbor changes propagate through the network with the delete_neighbors algorithm (Algorithm 5, in this paper’s appendix).
5 The Uniform SLN Protocol
After discussing the case of static SLN in the previous sections, we now model randomized P2P protocols. This section introduces a randomized, Continuous SLN called Uniform where each node chooses its height by picking a random real number in $(0,1)$ uniformly at random.
Definition 5.1 (Randomized Continuous SLN).
A randomized Continuous SLN is a probabilistic distribution over Continuous SLN.
Protocol 1 (Uniform).
Uniform is a randomized, continuous SLN that sets its heights to a real number in $(0,1)$ chosen uniformly at random.
The rest of this section presents the properties of Uniform with respect to routing path lengths and maximum degrees. We show that a Continuous SLN constructed with Uniform has an expected constant degree and an expected routing path length of $O(\log n)$. Also, we can rule out rare bad outcomes in practice: with high probability, the routing path length is $O(\log n)$ and the maximum degree is $O(\log n)$. Those results directly imply that the three primitive described in the previous section have a logarithmic message complexity with high probability. Interestingly, those results are quite general and hold for a vast family of continuous probability laws, namely, as soon as the probability that two heights are equal is zero.
Theorem 5.2 (Expected Path Length).
The routing path between any two nodes has length smaller than $2\ln n$ on expectation.
Proof 5.3.
@@ 5 The Uniform SLN Protocol
The last inequality holds as the harmonic sum until $n$, $1+1/2+\dots+1/n$, is lower than $1+\ln n$. The claimed upper bound on the expected path length follows as the same reasoning can be done for $X_{\text{down}}$.
Theorem 5.4 (Routing Path Length with High Probability).
The routing path between any two nodes has length smaller than $2e\cdot\ln n$ with probability greater than $1-\frac{2}{n}$.
Proof 5.5.
Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ . Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 5.2).
@@
$P[1_{w^{\prime}}=1\mid 1_{w}=1]$ $=P[h_{x}<h_{w^{\prime}}\forall x\in[w+1,w^{\prime}-1]\cap\max_{y\in[u,w]}h_{y}<h_{w}^{\prime}]$
$=P[h_{y}<h_{w}^{\prime}\forall y\in[u,w^{\prime}-1]]$
$=P[1_{w^{\prime}}=1]$
@@
$<\frac{2}{n}$
Theorem 5.6 (Expected Degree).
The expected degree of any node is lower than $4$.
Proof 5.7.
@@
$=4\cdot(1-\frac{1}{n/2+1})\leq 4$
Theorem 5.8 (Maximum Degree with High Probability).
The maximum degree is lower than $4\ln n$ with probability greater than $1-\frac{1}{n^{3}}$.
Proof 5.9.
Let $u$, $w$ and $w^{\prime}$ be three different nodes. We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to $1$ if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively. We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 5.4.
We first establish that each node taken individually has a degree lower than $\ln n$ with probability greater than $1-\frac{1}{n^{4}}$. Let $N_{u}$ be the random variable equal to the number of neighbors of $u$, we use the Chernov bound for Poisson trials used in the proof of Theorem 5.4:
@@
This last equality holds when $n$ is greater than a given constant, which is not a problem as we are interested in cases when $n$ is large. We conclude with a union bound: the probability that at least one node has a degree greater than $4\ln n$ is upper bounded by $\sum_{v\in V}P[N_{v}\geq 4\ln n]\leq\frac{1}{n^{3}}$ and the claim holds.
6 Learning-Augmented Continuous SLN

### v2

@@ 3 Optimum Static Skip List Network
OSSLN with Unbounded height (OSSLNU): We now present a dynamic programming algorithm for the OSSLN (4). Furthermore, we show that the algorithm’s complexity can be reduced to $\mathcal{O}(n^{3})$.
A crucial observation is that, for any solution to the OSSLNU problem, there exists a solution of cost less than or equal to that of the former, in which, $\forall[l,r],1\leq l\leq r\leq n$, there is a unique pivot node, i.e., a unique node whose height is the maximum height in that interval.
Lemma 2.
Let $h$ be an optimum solution to $\text{OSSLNU}(V,W)$. Then, there must exist a solution $h^{\prime}$ such that $Cost((V,h^{\prime},E^{\prime}),W)\leq Cost((V,h,E),W)$, where, for any interval $l,r],1\leq l\leq r\leq n,$ there is exactly one node $x\in[l,r]$ with height $h^{\prime}_{x}=\max_{w\in[l,r]}h^{\prime}_{w}$.
Proof.
@@
Since $\text{Cost}((V,h,E),W)=\sum_{(u,v)\in V}W(u,v)\cdot d(u,v)$ and $d(u,v)=|P_{h}(u,v)|$, and any path in $h^{\prime}$ is of at most the same size as in $h$, then, $\text{Cost}((V,h^{\prime},E^{\prime}),W)\leq\text{Cost}((V,h,E),W)$. Hence, the lemma holds. ∎
In OSSLNU, because we are only interested in the relative order of node heights, rather than in their absolute values, we have that $(nl=l)$ and $(nr=r)$, $\forall[l,r],1\leq l\leq r\leq n$, and the height parameter $h$ becomes unnecessary. This simplifies the recurrence relation, reducing the complexity to $\mathcal{O}(n^{3})$.
Theorem 1.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$. Let $\text{OSSLNU}(l,r,W)$ denote the minimum total cost contributed by all nodes $v\in[l,r]$. Let $x\in[l,r]$ be a candidate pivot node. Define $l$ as the leftmost index such that $\forall x,\in[l,r],\forall v\in[l,x],\,h_{v}\leq h_{x}$, and similarly define $r$ as the rightmost index such that $\forall v\in[x,r],\,h_{v}\leq h_{x}$. Then, the recurrence relation $\text{OSSNU}(l,r,W)=$
$\min_{x\in[l,r]}\Big(\text{OSSNU}(l,x-1,W)+\text{OSSNU}(x+1,r,W)+$
$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)\Big).$
Proof.
@@
Note that, by Lemma 2 once a pivot vertex $x\in[l,r]$ is chosen as the highest vertex within the interval $[l,r]$, we can assume, without loss of generality, that all vertices in the optimal solutions to the left subinterval [l,x1][l,x-1] and to the right subinterval $[x+1,r]$ have heights strictly lower than $h_{x}$.
@@
Base Case $(l=r)$: In this case, the interval contains a single vertex $x=l=r$. In this case, the cost reduces to the contribution of $x$ to the solution cost, which, from Lemma 1, and $L(x,h)=l$ and $R(x,h)=r$ as we demonstrated above, it follows that this expression corresponds exactly to:
@@
  • The cost of all paths that include $x$, which by Lemma 1 is given by the Equation 4.
Since we minimize this total over all pivot choices $x\in[l,r]$, the recurrence holds. ∎
OSSLN with Bounded Height: While OSSLNu optimizes the total routing cost, it offers no guarantees regarding individual node degrees or the average network degree. Consequently, a single node could theoretically accumulate a near-linear number of edges, rendering a physical network implementation infeasible.
Since the degree of a node is upper-bounded by twice its height, we can control the maximum degree by enforcing a height limit, $h_{\max}$. For the discrete SLN, capping the height at $h_{\max}$ strictly ensures that the maximum node degree does not exceed $2\times h_{\max}$.
Definition 8 (OSSLN with Bounded Height $h_{\max}$).
An OSSLN with bounded height $h_{\max}$ $\mathcal{N_{H}}=(V,h,E)$ is the discrete SLN (1) that minimizes the communication cost given by Equation 1 such that the height assignment for any of its nodes is at most $h_{\max}$.
In the OSSLN with bounded height, the height assignment is of the form $h:V\longrightarrow[1,h_{\max}]$, where $h_{\max}$ is a parameter, typically bounded by $O(\log{n})$ or $O(1)$.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$. Let $nl$, $1\leq nl\leq l$, denote the leftmost node id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost node id, such that $\forall v\in[r,nr],\,h_{v}\leq h$. Finally, let $\textsc{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all nodes $v\in[l,r]$.
Theorem 2.
If $x\in[l,r]$ is a candidate pivot node with height $h\in[1,h_{\max}]$, then the following recurrence relation holds $\forall(nl,l,x,r,nr)$, $1\leq nl\leq l\leq x\leq r\leq nr\leq n$:
$\min_{x\in[l,r]}\Bigg[\text{OSSLN}(l,x-1,nl,nr,h,W)+$
$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
Proof.
Base case $(l=r)$: In this case, we have that $(x=l)$ and the set of paths $x$ belongs to is given by Lemma 1 By setting $nl=L(x,h)$ and $nr=R(x,h)$ in the proof, the vertex contribution to the total cost is given by:
Let $1\leq l<r\leq n$ and $h\in[1,h_{\max}]$. Assume that the recurrence equation is true for all $h^{\prime},l^{\prime},r^{\prime}$ such that $(h^{\prime}<h)~\wedge~(r^{\prime}-l^{\prime})<(r-l)$. From previous step also assume that the equation is true for $(r^{\prime}-l^{\prime})=(r-l)$ with height $(h^{\prime}<h)$.
$\min_{x\in[l,r]}\Bigg[\textsc{OSSLN}(l,x-1,nl,nr,h,W)+$
$\textsc{OSSLN}(x+1,r,nl,nr,h,W)+$
and OSSLN(l,r,l,r,h1,W)OSSLN(l,r,l,r,h-1,W), which is precisely what we want to prove. ∎
The complexity of the above dynamic programming algorithm is $\mathcal{O}(n^{5}\cdot h_{\max})$, if we pre-compute an auxiliary matrix of aggregate weights $W^{\prime}(x,y)=\sum_{z\leq y}W(x,z),\forall(x,y),1\leq x\leq y\leq n$.
4 Continuous Skip-List Networks
This section introduces a generalization of SLN called C-SLN.
Definition 9 (Continuous Skip List Network (C-SLN)).
A C-SLN $\mathcal{N}=(V,h,E)$ is an SLN (1) except the height function $h$ is defined over the real numbers, i.e., $h:V\rightarrow\mathbb{R}$.
In this section, we show that C-SLN offer a set of properties similar to classic SLN with arguably simpler proofs. Next, we detail key procedures of C-SLN, namely join, leave and route.
4.1 Procedures
We first present the internal variables used in our procedures. In particular, a node $u$ holds the variables:
@@ 4.1 Procedures
The route algorithm (Procedure 1) takes the destination node id $v$ and hops from node to node until it finds the destination node, returning $(v,True)$ or reaches a dead end in some node $w\neq v$, return $(w,False)$.
// code running at node $u$
if $u$ = $v$ then
 return ($u$, True)
end if
else if $u<v$ then
 $S\leftarrow\{x\in N_{u}|u<x\leq v\}$
 if $S=\emptyset$ then
    return ($u$, False)
   end if
 else
    $w\leftarrow\max S$
    return $w.\text{route}(v)$
   end if
end if
else
 $S\leftarrow\{x\in N_{u}|v\leq x<u\}$
 if $S=\emptyset$ then
    return ($u$, False)
   end if
 else
    $w\leftarrow\min S$
    return $w.\text{route}(v)$
   end if
end if
Procedure 1 route($v$)
@@
The join algorithm (Procedure 2) inserts the node inside the Continuous SLN based on its anchor local variable. The potential neighbor changes of the other nodes propagate through the network with the find_neighbors algorithm (Procedure 4, in this paper’s appendix).
// code running at node $u$
if $\text{anchor}<u$ then
 $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
end if
else
 $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
end if
if $S=\emptyset$ then
 $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
   Update $D_{u}$
end if
else
 if $\text{anchor}<u$ then
    $\text{succ}\leftarrow\min S$
   end if
 else
    $\text{succ}\leftarrow\max S$
   end if
 $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
 $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{find\_neighbors}(u,h_{v},-\infty)$
   Update $D_{u}$
end if
Procedure 2 join()
@@
The leave algorithm (Procedure 3, in this paper’s appendix) shows how an already-inserted node can leave the network. The potential neighbor changes propagate through the network with the delete_neighbors algorithm (Procedure 5, in this paper’s appendix).
// code running at node $u$
if $\text{anchor}<u$ then
 $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
end if
else
 $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
end if
if $S=\emptyset$ then
 $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{delete\_node}(u,h_{v},-\infty)$
   Update $D_{u}$
end if
else
 if $\text{anchor}<u$ then
    $\text{succ}\leftarrow\min S$
   end if
 else
    $\text{succ}\leftarrow\max S$
   end if
 $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{delete\_node}(u,h_{v},-\infty)$
 $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{delete\_node}(u,h_{v},-\infty)$
   Update $D_{u}$
end if
Procedure 3 leave()
4.2 Uniform C-SLN
We now present a simple C-SLN called Uniform whose heights are chosen uniformly at random in $[0,1]$, and show that Uniform has the performances of state-of-the-art overlay network algorithms as both its routing search path and maximum degree are logarithmic with high probability. The properties of Uniform will be used to derive our main result on LASLiN in the next section.
Algorithm 1 (Uniform).
Uniform is a C-SLN whose initialization procedure sets its heights to a real number in $[0,1]$ chosen uniformly at random.
Theorem 3 (Expected Routing Path Length).
For Uniform, the routing path between any two nodes has length smaller than $2\ln n$ on expectation.
Proof.
@@ 4.2 Uniform C-SLN
The last inequality holds as the harmonic sum until $n$, $1+1/2+\dots+1/n$, is lower than $1+\ln n$. The claimed upper bound on the expected path length follows as the same reasoning can be done for $X_{\text{down}}$. ∎
Theorem 4 (Routing Path Length with High Probability).
For Uniform, the routing path between any two nodes has length smaller than $2e\cdot\ln n$ with probability greater than $1-\frac{2}{n}$.
Proof.
Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ . Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 3).
@@
$P[1_{w^{\prime}}=1\mid 1_{w}=1]$ $=$
$P[h_{x}<h_{w^{\prime}}\forall x\in[w+1,w^{\prime}-1]\cap\max_{y\in[u,w]}h_{y}<h_{w}^{\prime}]$ $=$
$P[h_{y}<h_{w}^{\prime}\forall y\in[u,w^{\prime}-1]]$ $=$
$P[1_{w^{\prime}}=1]$
@@
$<\frac{2}{n}\qed$
Theorem 5 (Expected Degree).
For Uniform, the expected degree of any node is lower than $4$.
Proof.
@@
$=4\cdot(1-\frac{1}{n/2+1})\leq 4\qed$
Theorem 6 (Maximum Degree with High Probability).
For Uniform, the maximum degree is lower than $4\ln n$ with probability greater than $1-\frac{1}{n^{3}}$.
Proof.
Let $u$, $w$ and $w^{\prime}$ be three different nodes. We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to $1$ if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively. We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 4.
We first establish that each node taken individually has a degree lower than $\ln n$ with probability greater than $1-\frac{1}{n^{4}}$. Let $N_{u}$ be the random variable equal to the number of neighbors of $u$, we use the Chernoff bound for Poisson trials used in the proof of Theorem 4:
@@
This last equality holds when $n$ is greater than a given constant, which is not a problem as we are interested in cases when $n$ is large. We conclude with a union bound: the probability that at least one node has a degree greater than $4\ln n$ is upper bounded by $\sum_{v\in V}P[N_{v}\geq 4\ln n]\leq\frac{1}{n^{3}}$ and the claim holds. ∎
5 Learning-Augmented C-SLN